OpenStudy (anonymous):

complete the square to rewrite the quation in vertex form. a) y= -x^2+2x-7 b) y= 2x^2+6x-5 (they are 2 separate problems) please help me. thanks so much.

OpenStudy (anonymous):

a) y=-(x^2-2x+7)=-[(x-1)^2+6]=-(x-1)^2-6 b)y=2(x^2+3x-5/2)=2[(x+3/2)^2-19/4]=2(x+3/2)^2-19/2